n= 1000, n= 60, mean = 75, standard deviation= 6, probability= p(x<76.5)


1
Expert's answer
2021-04-29T17:32:37-0400

Here,

"Mean\ \mu=75\\Standard\ Deviation\ \sigma=6\\\\N=1000\\n=60"


"z=\dfrac{x-\mu}{\sigma}=\dfrac{76.5-75}{6}=0.25"


So, "P(x<76.5)=P(z<0.25)=0.5987"